### v1

# Distribution Transformers: Fast Approximate Bayesian Inference With On-The-Fly Prior Adaptation

Abstract
While Bayesian inference provides a principled framework for reasoning under uncertainty, its widespread adoption is limited by the intractability of exact posterior computation, necessitating the use of approximate inference. However, existing methods are often computationally expensive, or demand costly retraining when priors change, limiting their utility, particularly in sequential inference problems such as real-time sensor fusion. To address these challenges, we introduce the Distribution Transformer—a novel architecture that can learn arbitrary distribution-to-distribution mappings. Our method can be trained to map a prior to the corresponding posterior, conditioned on some dataset—thus performing approximate Bayesian inference. Our novel architecture represents a prior distribution as a (universally-approximating) Gaussian Mixture Model (GMM), and transforms it into a GMM representation of the posterior. The components of the GMM attend to each other via self-attention, and to the datapoints via cross-attention. We demonstrate that Distribution Transformers both maintain flexibility to vary the prior, and significantly reduces computation times—from minutes to milliseconds—while achieving log-likelihood performance on par with or superior to existing approximate inference methods across tasks such as sequential inference, quantum system parameter inference, and Gaussian Process predictive posterior inference with hyperpriors.
Machine Learning, ICML, Approximate Inference, Transformers
1 Introduction
Modern machine learning has transformed many data-rich domains (LeCun et al., 2015), yet critical applications like healthcare diagnostics and autonomous vehicle safety demand more than just predictive accuracy (Hüllermeier & Waegeman, 2019). These high-stakes settings, often characterized by limited or noisy data, require robust uncertainty estimation and principled integration of domain knowledge (Gawlikowski et al., 2021)—capabilities naturally addressed through probabilistic reasoning.
Bayesian inference is a cornerstone of probabilistic machine learning, encoding prior beliefs and providing the tools to incorporate new observations while providing principled uncertainty quantification; essential for critical decision-making under uncertainty. However, Bayesian inference is typically intractable, limiting its real-world adoption. To counteract this problem, the field of approximate inference has emerged, providing a range of methods: Variational inference (VI) methods (Blei et al., 2016) can approximate posterior densities but require expensive optimization procedures at inference time (as opposed to training time, relevant for later methods). While Markov Chain Monte Carlo (MCMC) (Christian & Casella, 2011) methods provide asymptotically-exact posterior samples, they neither directly estimate the density nor offer the rapid inference often required in practical applications.
Recent advances in deep learning have opened new avenues for Bayesian inference. Prior Fitted Networks (PFNs) (Muller et al., 2021) demonstrated that transformer (Vaswani et al., 2017) architectures can perform Bayesian inference through a single forward pass, dramatically reducing inference time compared to traditional methods. However, PFNs are limited by their dependence on a fixed prior, requiring costly retraining for each new prior distribution encountered. In practice, it is often desirable, and in some applications essential, to be able to change the prior rapidly at inference time.
While these methods have their strengths, there remains a clear need for approximate inference methods that can provide accurate posterior density estimates with minimal computational overhead while maintaining the flexibility to modify priors at inference time. To address this need, we introduce Distribution Transformers (DTs), a novel transformer-based architecture that combines the flexibility of VI with the efficiency of PFNs. Our method generates posterior density estimates through a single forward pass while handling a range of prior distributions from a given parametric family without retraining. Unlike traditional VI approaches that often rely on simple parametric families, DTs leverage a expressive distributional family that can universally approximate smooth densities. Figure 2 illustrates the architectural design of DTs. Fundamentally, DTs approximate the prior with a Gaussian Mixture Model (GMM) represented as an embedded sequence in the latent space of the transformer, then condition this GMM on observations using a permutation-equivariant transformer decoder, yielding a GMM approximation to the posterior.
DTs are trained using entirely sample-based supervised learning by defining a joint distribution over the prior, parametrized by $\phi$, the latent variable $x$, and the observations $z$. We demonstrate that this training approach yields an approximator of exact posterior prediction. This combination of speed, flexibility, and expressiveness enables accurate Bayesian inference in scenarios where existing methods face fundamental limitations. Concretely, DTs address the key challenges discussed above, providing accurate posterior density estimates for a wide range of priors. Like PFNs, our approach only requires a way to sample from the chosen prior family and an arbitrary likelihood model, and is widely applicable without significant task-specific analysis. We make the following contributions:
  1. 1.
    We introduce Distribution Transformers (DTs)—a novel architecture that leverages the power of universally-approximating Gaussian Mixture Models and Transformers to express arbitrary distribution-to-distribution mappings.
  2. 2.
    We show that DTs can be trained to perform more accurate approximate Bayesian inference, and up to orders of magnitude faster, than existing approximate inference methods, while also allowing on-the-fly updates to the prior at inference time without retraining, a never before accomplished feat.
  3. 3.
    We empirically demonstrate near-exact inference capabilities on challenging problems, including joint PPD-hyperposterior approximations for Gaussian Processes (GPs), parameter inference for quantum systems and a real-world sequential inference task, while remaining fast enough for real-time applications.
1.1 Related Work
[…]
Variational inference (VI) aims to approximate the posterior density through fitting a density of a particular family (typically a Gaussian) to the posterior by maximizing a variational lower bound on the evidence. Notable modern extensions include stochastic variational inference (SVI) (Hoffman et al., 2013), which utilizes backpropagation to optimize a stochastic approximation of the variational objective, black-box variational inference (Ranganath et al., 2014), which provides widely-applicable, model-agnostic tools for optimization of this variational objective, and amortized variational inference (AVI) (Kingma & Welling, 2014; Ganguly et al., 2023), which efficiently solves many variational inference problems in parallel through the learning of an inference network mapping from observations to approximate posteriors. These methods often provide acceptable estimates for the posterior density, but necessarily involve an expensive optimization routine at inference time, are not asymptotically exact, and often underestimate the posterior variance. Nonetheless, VI remains the most popular approximate inference technique in use today, and we will use a variant as a baseline in our empirical studies.
Recent neural approaches to approximate inference leverage the expressivity of deep learning architectures while aiming to maintain theoretical guarantees. Prior Fitted Networks (PFNs) (Muller et al., 2021) learn to perform approximate inference through sample-based supervised learning without the need for posterior densities (which are instead accessed via samples from the joint distribution over $x$ and $z$), enabling inference in a single forward pass. They provide rapid inference for changes in dataset, provide a posterior density estimate, and allow for almost arbitrary priors, but require complete retraining upon changes to the prior distribution, and use a Riemann Distribution, which can struggle to model smooth, heavy-tailed posteriors. As a close competitor to our method, we will also adopt a PFN baseline for our posterior studies.
Continuous normalizing flows (Rezende & Mohamed, 2015) learn a mapping from samples of some simple distribution to samples of the target posterior density, leveraging automatic differentiation to provide a flexible, closed-form density using a change of measure. However, the mapping they learn is fixed, meaning the prior is set at training time, and observations can only be incorporated via sufficient statistics of the simple distribution, limiting their suitability for our problem setting. Moreover, they must be constructed using only fully reversible components, which can limit expressivity. Classical approaches include the Laplace approximation (Tierney & Kadane, 1986) and expectation propagation (EP) (Minka, 2001). The Laplace approximation fits a Gaussian distribution based on the local geometry at the posterior mode, but lacks invariance to reparametrization. EP iteratively refines an isotropic Gaussian approximation to the posterior, but struggles with numerical stability.
Sampling-based methods, most notably Markov chain Monte Carlo (MCMC) methods, provide samples from the posterior which can be used to compute stochastic estimates for expectations with respect to said posterior. These are primarily based on access to the unnormalized posterior via the numerator of Bayes’ theorem. When a differentiable prior and likelihood model are available, the state-of-the-art No U-Turn Sampler (NUTS) (Hoffman & Gelman, 2011) can be used. NUTS is based on Hamiltonian Monte Carlo (HMC) (Brooks et al., 2012), which provides more efficient generation of posterior samples without problem-specific tuning of the algorithm. MCMC methods are asymptotically exact and almost universally applicable, but are typically expensive due to random walk behavior. They also do not provide direct access to a posterior density, and empirical density estimates can be sample inefficient.
Bayesian quadrature-based methods (Osborne et al., 2012) assign a probabilistic model, typically a GP, to the integrand of the evidence $p(z)$ (and in some cases the predictive posterior distribution PPD $p(y|x,z)=\int p(y|x)p(x\mathop{|}z)$), and use probabilistic numerical integration to estimate the evidence and PPD directly. These methods provide accurate uncertainty measures, and can handle noisy, black-box likelihood models, but are typically computationally expensive and scale poorly to high-dimensional problems, although recent advances in sparse-GPs (Binois & Wycoff, 2021) provide promising avenues to alleviate these scaling issues.
1.1.1 Sequential Inference
Sequential inference over time series requires dedicated methods, referred to as Bayesian filtering (when causality is required) or smoothing (when not). The extended Kalman filter (EKF) linearizes dynamics and observation models, then approximates the distribution over the latent state as a closed-form Gaussian. With modern automatic differentiation, this method is widely applicable and computationally efficient, but struggles in practice due to the overly restrictive assumptions induced by linearization and Gaussian distributions. Particle filters (PFs), or sequential Monte Carlo methods (Wills & Schön, 2023), maintain a population of particles which are distributed according to the posterior distribution each time step, but can be expensive and only provide an empirical density. The unscented Kalman filter (Julier & Uhlmann, 1997) maintains a minimal population of particles so is computationally efficient, but faces issues of numerical stability and also encodes a Gaussian assumption thus cannot handle multimodality. We compare our method against an EKF and a PF baseline.
2 Preliminaries
2.1 Transformers
The transformer architecture (Vaswani et al., 2017) has revolutionized deep learning, achieving state-of-the-art performance across domains including language modelling (Brown et al., 2020), computer vision (Dosovitskiy et al., 2020), and Bayesian inference (Muller et al., 2021). At its core, transformers learn mappings between sequences of tokens through the attention mechanism (Bahdanau et al., 2014), which enables parallelized information flow between sequence elements. This mechanism, combined with token-wise MLP layers, creates a parameter-efficient architecture capable of processing sequence elements in parallel.
Two theoretical properties of transformers are central to our work. First, transformers are universal sequence-to-sequence approximators (Yun et al., 2019), capable of representing arbitrary mappings between sequences. Second, in the absence of positional encodings, transformers are permutation equivariant with respect to the input sequence—a property we will exploit in Section 3.
Our method specifically employs the transformer decoder architecture. This architecture extends the base transformer by incorporating global cross-attention layers, allowing each position in the input sequence to attend to a separate context sequence. This cross-attention mechanism provides a natural framework for conditioning sequence transformations on observed data.
2.2 Gaussian Mixture Models
Gaussian Mixture Models (GMMs) are flexible probability distributions whose density is a weighted sum of Gaussian components, i.e. $q(x)=\sum_{i}w_{i}\mathcal{N}(x;\mu_{i},\Sigma_{i})$ where $\mathcal{N}(x;\mu,\Sigma)$ is a Gaussian density over $x,w_{i}\in[0,1]$, $\sum_{i}w_{i}=1$, $\boldsymbol{\mu}_{i}\in\mathbb{R}^{n}$, and $\boldsymbol{\Sigma}_{i}\in\mathbb{S}_{++}^{n}$, illustrated in Figure 1. While GMMs are widely used in clustering (Dempster et al., 1977), and latent variable modelling (Bishop, 2006), we focus on their role as universal approximators of smooth probability distributions (Goodfellow et al., 2016; Calcaterra & Boldt, 2008)—a property we exploit in Section 3. This universality extends to distributions on compact domains under appropriate change of measure, also illustrated in Figure 1. While the idea of using GMMs for function approximation is not new to deep learning (for example Bishop (1994) proposes the use of a GMM to model uncertainty in the output of a neural network), the idea of operating end-to-end on a distribution represented as a GMM is novel.
Figure 1: Various log-warped GMM approximation to an inverse-gamma prior distributions. Note that even with only five GMM components, the approximation is visually almost indistinguishable from the target distribution. This is true for many frequently encountered distributions in Bayesian inference.
A key property of GMMs is their natural representation as an unordered sequence of component parameters. Specifically, a $k$-component GMM over $\mathbb{R}^{n}$ is parametrized by $\boldsymbol{\theta}=\{(w_{i},\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})\}_{i=1}^{k}$. This representation is permutation invariant—the ordering of components does not affect the resulting distribution.
Fitting a GMM to a given probability distribution is non-trivial, with methods such as expectation maximization and variational methods suffering from similar problems to their approximate inference counterparts. We will show that DTs naturally provide a mapping from the parameters of a given distribution to an approximating GMM without introduction of additional model parameters.
2.3 Amortization
Amortized methods have received significant attention from the machine learning, statistics, and optimization communities, owing to their ability to address computational bottlenecks in repetitive tasks. These methods learn to solve a class of optimization or simulation problems by incurring a substantial up-front computational cost during training, assuming the existence of some common structure between the solutions to these problems. This initial investment is offset by enabling the rapid solution of many subsequent problems, significantly reducing the overall computational burden over the model’s lifetime.
Amortized variational inference (AVI) (Ganguly et al., 2023; Kingma & Welling, 2014) exemplifies this paradigm, where a neural network is trained to approximate posterior distributions across multiple datasets, bypassing the need for iterative optimization for each new dataset. Similarly, PFNs (Muller et al., 2021) implicitly amortize inference through the learning of a mapping from observations to approximate posterior distributions without iterative updates. These examples highlight how amortization enables scalability and efficiency in otherwise computationally intensive workflows. As we shall show in Section 3, our work is also well-set in the context of amortized methods.
3 Distribution Transformers
Given a prior distribution $p(x\mathop{|}\phi)$ from a parametric family with parameters $\phi\in\Phi$ and observations $z\in\mathcal{Z}$ governed by likelihood $p(z\mathop{|}x)$, Bayesian inference aims to compute the posterior $p(x\mathop{|}z,\phi)$. Amortized approximate Bayesian inference reframes this as learning a mapping $\Phi\times\mathcal{Z}\to\mathcal{Q}$, where $\mathcal{Q}$ is a space of approximate posteriors. We introduce the Distribution Transformer (DT), a transformer-based architecture that directly maps priors and observations to posteriors. A core challenge in Bayesian inference is representing arbitrary probability distributions in a form suitable for neural networks. Our first key innovation is to represent all distributions as Gaussian Mixture Models (GMMs), which approximate any continuous density arbitrarily well. Our second key innovation is a transformer architecture that processes these mixtures as unordered sequences, preserving probabilistic structure while enabling expressive, scalable inference.
Figure 2: Architecture diagram for a distribution transformer. Observations, e.g. from a dataset or a sensor measurement, are transformed to a set of tokens in the latent space via a distinct learnable embedding for each datasource. Priors are represented as a set of embedded GMM components in the latent space via a learnable embedding acting on their parameters. The distribution transformer itself, a transformer decoder, learns to map the prior to the posterior in the latent space, incorporating information from the embedded observations via cross attention. A learnable unembedding acts token-wise on both the prior and posterior latent GMM representations to give a GMM approximation for both the prior and posterior distributions, with which we estimate our loss function $\ell_{\theta}^{\prime}.$
Figure 2 illustrates the DT architecture in detail. Conceptually, the DT architecture can be broken into four parts: the prior embedding, observation embeddings, transformer decoder and GMM unembedding.
Obtaining a GMM representation of an arbitrary prior is nontrivial. Inspired by Bishop (1994), we introduce a learnable embedding network that maps a prior’s parameters to a length-$k$ unordered sequence in the transformer’s latent space, representing an embedding of a $k$-component GMM approximation to the prior.
Observations $z$ vary widely (e.g., sensor readings, datasets). We use learnable embeddings tailored to different data sources, and embed datasets as sequences of data-label pairs embedded token-wise by a single embedding model. If a predictive posterior is needed, the query point is embedded separately. Observations are combined a unified latent sequence.
The transformer decoder maps the latent GMM prior representation to a posterior representation, conditioned on the observations via global cross-attention. We omit positional encodings to preserve permutation equivariance, aligning with the permutation invariance of the GMM sequence representation.
A GMM posterior approximation is then obtained via a learnable unembedding that acts component-wise on the posterior latent GMM unordered sequence, producing logits and normal densities. A cross-sequence softmax converts logits into component weights, and the approximating GMM can be constructed through summation of these components, achieving end-to-end permutation invariance of the architecture, as required.
Now that we have an architecture capable of mapping between distributions, we propose a sample-based training scheme with which to train our architecture to perform Bayesian inference. We must first introduce the concept of meta-priors $p(\phi)$—priors over priors representing the expected distribution of priors encountered by the algorithm. The only constraint on these meta-priors is that they can be sampled from, and can otherwise be quite complicated. For instance, in vehicle tracking, a meta-prior could constrain Gaussian means to a city’s road network while shaping covariance to reflect realistic uncertainties. Using this meta-prior, we can specify the joint distribution $p(\phi,x,z)$ hierarchically as $p(\phi)p(x\mathop{|}\phi)p(z\mathop{|}x)$. We may also specify a mapping $f(\cdot)$ from the sample space of interest to the sample space of the approximating GMM $\mathbb{R}^{n}$, for example to account for priors with finite support, essentially specifying a change of measure ensuring the probabilistic properties of the approximation are maintained. In this case, we denote the GMM itself as $q_{\theta}(f(x)\mathop{|}z,\phi)$, inducing the warped GMM $q_{\theta}(x\mathop{|}z,\phi)\approx p(x\mathop{|}z,\phi)$ under change of measure.
Algorithm 1 Training a Distribution Transformer
  Input: A joint distribution $p(\phi,x,z)$ over priors, latent variables and observations, the number of training iterations $m$, the batch size $b$, the number of GMM components $k$, and a mapping $f$ from the sample space of interest to the sample space of the GMM.
  Output: A mapping from $\phi$ and $z$ to warped GMMs $q_{\theta}(x\mathop{|}\phi)$ and $q_{\theta}(x\mathop{|}z,\phi)$ approximating the prior $p(x\mathop{|}\phi)$ and posterior $p(x\mathop{|}z,\phi)$ respectively.
  for $i=1$ to $m$ do
     Sample $\phi_{i}$, $x_{i}$ and $z_{i}$ from $p(\phi,x,z)$ for $i=1:b$;
     Estimate loss ^$\hat{\ell}^{\prime}_{\theta}=-\sum_{i=1}^{b}\log q_{\theta}(f(x_{i})\mathop{|}\phi_{i})+\log q_{\theta}(f(x_{i})\mathop{|}z_{i},\phi_{i})$;
     Update DT parameters with gradient descent on $\nabla_{\theta}\hat{\ell}^{\prime}_{\theta}$;
  end for
Outlined in Algorithm 1, DTs are trained to minimize $\ell_{\theta}^{\prime}$. Using meta-priors and a sample-space transform $f$, we extend the loss function proposed by Muller et al. (2021) another meta-level, defined as $\ell_{\theta}=\mathbb{E}_{p(\phi,x,z)}\left[-\log q_{\theta}(f(x)\mathop{|}z,\phi)\right]$. We show that this is equivalent to direct minimization of the KL-Divergence between the true posterior $p(x\mathop{|}z,\phi)$ and the GMM approximation $q_{\theta}(x\mathop{|}z,\phi)$:
Proposition 3.1.
The proposed loss $l_{\theta}$ equals the expected KL-Divergence $\mathbb{E}_{p(\phi,z)}\left[\text{KL}\left[p(\cdot\mathop{|}z,\phi)\mathop{||}q_{\theta}(\cdot\mathop{|}z,\phi)\right]\right]$ between $p(\cdot\mathop{|}z,\phi)$ and $q_{\theta}(\cdot\mathop{|}z,\phi)$ up to an additive constant. Proof in Appendix A.
The loss $l_{\theta}$ can be calculated stochastically using samples from the joint distribution $p(\phi,x,z)$, alleviating any need to directly access or sample from the posterior density.
Finally, DTs can jointly approximate priors and posteriors without additional model parameters. Applying the unembedding to the prior sequence yields a GMM approximation $q_{\theta}(x\mathop{|}\phi)$, extending DTs to mappings $\Phi\times\mathcal{Z}\rightarrow\Theta\times\Theta$. To ensure consistency, we introduce a prior loss:
$\ell_{\theta}^{\text{prior}}=\mathbb{E}_{p(\phi,x)}\left[-\log q_{\theta}(x\mathop{|}\phi)\right],$
leading to the combined objective: $\ell_{\theta}^{\prime}=\ell_{\theta}^{\text{prior}}+\ell_{\theta}$
4 Empirical Studies
We study the behaviour of our method in three settings: approximation of a tractable posterior, approximation of intractable posteriors, and a real-world sensor fusion problem posed as sequential inference. In the prior two settings, we benchmark our method against SVI (Hoffman et al., 2013), implemented in PyTorch (Paszke et al., 2019) with GPU parallelization, and PFNs (Muller et al., 2021), also implemented in PyTorch, fitting a Riemann distribution with the same number of model outputs as our DT. In the absence of a fixed prior, we train PFNs using the same sampling scheme as our method, effectively marginalizing out the meta-prior, leaving a less informative prior. We expect PFNs to perform well when the meta-prior is narrow and poorly when wide. For the latter experiment, we benchmark against the industry standard extended Kalman filter (EKF), and a particle filter (PF), again with GPU implementation. For the PF, we obtain a density via Gaussian kernel density estimation.
4.1 Analytical Verification Study
In specific cases, where the adopted prior is of a conjugate family to the likelihood, the posterior is tractable. We first verify that our approach does indeed perform approximate inference, for the case of an inverse-gamma prior and normal-variance likelihood. We choose a meta-prior consisting of independent inverse-gamma distributions over the rate and shape parameters, and test our approach on both narrow and wide parameter settings.
Figure 3: Ground truth, PFN, and 2 and 5 component DT posterior densities for an inverse-gamma prior with an (a) narrow and (b) wide meta-prior. Both variants of the DT fit the true posterior well, in both cases with the 5 component DT almost indistinguishable from the ground truth. The PFN’s shape is correct in both cases, and fits the ground truth correctly (up to the limits of the Riemann distribution) for the narrow meta-prior, but as expected completely fails to fit the ground truth for the wide meta-prior, given the lack of prior. In any case, for a given number of model outputs the DT provides a much tighter fit to the ground distribution.
Table 1: Results for Experiment 4.1 (best), tested on and timed over 1000 sampled unseen problems. Expected KL-Divergences with the true posterior (along with 95% confidence intervals) are given for the narrow (above) and wide (below) meta-priors. DT-$k$ refers to a $k$-component Distribution Transformer. Note first that both variants of our proposal achieve better posterior KL-Divergences than an equivalent PFN and SVI for both meta-prior settings, while performing inference orders of magnitude faster than SVI. This difference is particularly apparent for the wide meta-prior, where the PFN fails to fit the posterior entirely.
Method KL-Divergence
Inference Time per
1000 Problems (s)
SVI
0.0423$\pm$0.0003
0.0558$\pm$0.0016
148
PFN
0.814±$\pm 0.801^{*}$
726±$\pm 1070^{*}$
0.0144
DT-2
0.0045$\pm$0.0001
0.0058$\pm$0.0002
0.0150
DT-5
0.0004$\pm$0.0000
0.0003$\pm$0.0000
0.0156
Figure 3 demonstrates that for both wide and narrow meta-priors, the Distribution Transformer does indeed learn to perform Bayesian inference and provides excellent approximations to the posterior, even under change of prior. Figure 1 shows the GMM approximations provided by the DT for this study, demonstrating that even with no additional model parameters a high-quality mapping to a GMM approximation of the prior is achieved. It is clear that while PFNs perform as well as the Riemann distribution allows for the fixed-prior case, as expected in the case of variable priors they fail to fit the posterior at all. This is further demonstrated in Table 1, which shows that not only do we achieve speeds close to that of PFNs, which are slightly faster due to the lighter-weight architecture, but even significantly outperform the much slower SVI in terms of posterior KL-Divergence. These performance gains can be attributed to the expressive GMM adopted by our method, and the efficient transformer-based architecture. An interesting observation, is the extremely high uncertainty in the estimate for the PFN posterior KL-Divergence, marked *. This can be attributed to a failing of the Riemann distribution in this setting—the half-Gaussian tail adopted by the Riemann distribution has variance fitted to the (marginal) prior, meaning for certain observations (or priors), the true posterior has significant probability mass in the right tail which the Riemann distribution cannot express, leading to a skewed distribution for the expected KL-Divergence with a misleading 95%-confidence interval.
4.2 Posterior Approximation Studies
We now move our attention to problems where the posterior is intractable, as is more often the case.
4.2.1 Gaussian Process Joint Predictive Posterior and Hyperposterior
When modelling data with a Gaussian Process, it is common to assign priors to hyerparameters, known as hyperpriors. These hyperpriors render the predictive posterior intractable, and moreover, the posterior for the hyperparameters, or hyperposterior, is also intractable. Existing techniques tackle these distributions separately, and are plagued by the aforementioned issues. We now demonstrate that our method can quickly perform approximate inference jointly over both the predictive posterior and hyperposterior.
In Table 2 we show that we outperform existing methods in terms of NLL for both PPD and hyperposterior, as expected, while also being the fastest method in terms of runtime. In Figure 4, we show an example PPD for for one set of datapoints and are clearly closer to the oracle PPD that has the knowledge of the true lengthscale.
Figure 4: Example plot for the GP predictive experiment with hyperpriors, using 10 GMM components. Here we put an InverseGamma(1,2) prior on the lengthscale. We plot our model’s predictive posterior in blue and PFNs in green. In orange, we show the oracle that is the exact GP fit with the true lengthscale value (which is unobserved for the other methods). We see that PFNs overestimate the confidence intervals due to the fact that they do not take the prior, particularly that of the lengthscale, into account.The Riemann distribution of the PFN uses 30 buckets.
Table 2: Expected NLL for the marginal PPD and the marginal hyperposterior (denoted Hyper), and inference time per 1000 problems (denoted Runtime), for Experiment 4.2.1. VI is not evaluated for the PPD, as is convention. As expected, we outperform in both categories. Note that PFNs suffer the same issue with the hyperposterior NLL confidence here as the one in Experiment 4.1.
Method Expected NLL
PPD Hyper Runtime (s)
VI 1.97 $\pm$ 0.53 131
PFN 0.00 $\pm$ 0.10 3.21 $\pm$ 4.40 1.03
DT -0.32 $\pm$ 0.12 0.81 $\pm$ 0.08 0.96
4.2.2 Quantum System Parameter Inference
An interesting example of an inference problem involving genuine randomness with real-world implications is parameter inference for a quantum system. For this experiment, we infer the unknown parameter $\Delta$ for a two-level quantum system with Hamiltonian $H=\Delta\sigma_{x}+(1-\Delta)\sigma_{z}$, where $\sigma_{x}$ and $\sigma_{z}$ are the Pauli X and Z matrices respectively. We model observations as 10 independent experiment runs, subject to uncertainty in initial state preparation and measurement times, and generated via a GPU-implemented simulation.
Figure 5: Expected negative log-likelihood against batch inference time for SVI, a PFN and a 5-component DT. Note that even given orders of magnitude more computation time, SVI cannot match the performance of our method, again demonstrating the power of our GMM approximation. Note that this problem is particularly challenging for VI, as the likelihood must be marginalized with respect to the uncertainty in the initial state and measurement time, which is not tractable and must be estimated stochastically, increasing the time per iteration.
Figure 5 illustrates the power of our approach, achieving better log-likelihood performance (and therefore a smaller KL-Divergence with the true posterior, indicating a better fit), than both competitor methods. This is a problem setting where variable priors are important, as they can be subjective in this setting, and prior sensitivity analysis can be important. To enable such an analysis, fast inference is also important. This illustrates the practical advantage of our approach over PFNs and VI well, as we explicitly target these problems.
4.3 Sequential Inference Study
The industry standard for real-time sequential inference (or Bayesian filtering), and therefore sensor fusion, is the Extended Kalman Filter (EKF). The EKF linearizes system dynamics and observation models, and approximates all sources of uncertainty as Gaussian. However, these assumptions are rarely reflected in reality. We tackle a sensor fusion problem consisting of 2-dimensional linear dynamics, modelled as a 4-dimensional state space (displacement and velocity), with indirect measurements of displacement provided by two independent, non-linear, non-Gaussian, sensors. Our approach is likely to be successful here, as prior time spent training is almost irrelevant, and fast inference with variable priors is the priority. We also validate against a particle filter, which handles these non-ideal conditions well at the expense of computation time.
Figure 6: Marginal filter densities of horizontal displacement for both an EKF and a 4-component Distribution Transformer. Clearly, the DT tracks the true trajectory much more accurately, which is reflected in the superior expected NLL reported in Table 3. Note that the DT generally has a higher uncertainty than the EKF, demonstrating proper handling of the complex uncertainty structure of the observations.
Table 3: Expected NLL and iteration time (prediction, update and distribution generation) for EKF, PF, and DT. Note that our method vastly outperforms the EKF in terms of NLL with a small cost in iteration time, while almost matches the close to ground-truth PF’s NLL while achieving close to 50$\times$ speedup.
Method Expected NLL
Iteration Time for
100 Series Batch (s)
EKF 95.9$\pm$4.40 0.010
PF -0.244$\pm$0.047 0.818
DT -0.197$\pm$0.040 0.017
Figure 6 clearly demonstrates that our approach tracks the true state well, while the EKF fails to track the true state at all. Table 3 confirms this, and shows that our approach sacrifices little in terms of iteration time, which upper bounds the frequency at which observations can be processed in real time, compared to the EKF. As expected, the close-to-ground truth PF achieves marginally better NLL than our method, at the expense of a significant slowdown.
5 Conclusion
In summary, Distribution Transformers (DTs) introduce a powerful framework for approximate Bayesian inference by combining universally-approximating Gaussian Mixture Models with Transformer architectures. Our empirical results demonstrate that DTs not only achieve superior inference accuracy and speed compared to existing methods, but also enable dynamic prior updates without retraining—a unique capability in the field. Furthermore, DTs show near-exact inference performance on challenging tasks including GP hyperposterior estimation, quantum parameter inference, and real-world sequential inference, while maintaining the computational efficiency needed for practical applications.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Proof of Proposition 3.1
Proof.
This equality can shown with a simple derivation:
$\ell_{\theta}$ $=\mathbb{E}_{p(\phi,x,z)}\left[-\log q_{\theta}(f(x)\mathop{|}z,\phi)\right]$
$=\mathbb{E}_{p(\phi,z)}\left[\mathbb{E}_{p(x\mathop{|}\phi,z)}\left[-\log q_{\theta}(f(x)\mathop{|}z,\phi)\right]\right]$
$=\mathbb{E}_{p(\phi,z)}\left[\mathbb{E}_{p(x\mathop{|}\phi,z)}\left[-\log q_{\theta}(x\mathop{|}z,\phi)+\log\mathop{|}\det J_{f}(x)\mathop{|}\right]\right]$
$=\mathbb{E}_{p(\phi,z)}\left[\mathbb{E}_{p(x\mathop{|}\phi,z)}\left[\log p(x\mathop{|}z,\phi)-\log p(x\mathop{|}z,\phi)-\log q_{\theta}(x\mathop{|}z,\phi)+\log\mathop{|}\det J_{f}(x)\mathop{|}\right]\right]$
$=\mathbb{E}_{p(\phi,z)}\left[\mathbb{E}_{p(x\mathop{|}\phi,z)}\left[\log\frac{p(x\mathop{|}z,\phi)}{q_{\theta}(x\mathop{|}z,\phi)}\right]\right]+\mathbb{E}_{p(\phi,z)}\left[\mathbb{E}_{p(x\mathop{|}\phi,z)}\left[\log\mathop{|}\det J_{f}(x)\mathop{|}-\log p(x\mathop{|}z,\phi)\right]\right]$
$=\mathbb{E}_{p(\phi,z)}\left[\text{KL}\left[p(x\mathop{|}z,\phi)\mathop{||}q_{\theta}(x\mathop{|}z,\phi)\right]\right]+\mathbb{E}_{p(\phi,x,z)}\left[\log\mathop{|}\det J_{f}(x)\mathop{|}-\log p(x\mathop{|}z,\phi)\right],$
where $J_{f}(x)$ denotes the Jacobian of $f(\cdot)$ evaluated at $x$, the third line follows from change of measure, and the term $\mathbb{E}_{p(\phi,x,z)}\left[\log\mathop{|}\det J_{f}(x)\mathop{|}-\log p(x\mathop{|}z,\phi)\right]$ is constant with respect to $\theta$. ∎
Appendix B General Architecture Details
Learnable prior embeddings typically consist of a multi-layer perceptron (MLP) with one hidden layer, usually of half the size of the transformer latent space. For GMM priors, the prior embedding acts elementwise and consists of a logarithm applied to the component weight, a Cholesky decomposition then logarithm of diagonal elements applied to the covariance matrix, and a flattening into a vector before passing through the MLP straight to the latent unordered sequence. For general priors, simple transformations are applied to parameters, for example positive parameters are passed through a logarithm, before being passed through the MLP to a single vector in the latent space. Then, distinct MLPs act on this vector to yield the components of the latent unordered sequence.
Learnable observation embeddings always consist only of an MLP, but theoretically could be extended to include inductive biases suitable for the nature of each observation. These embeddings are not processed further, and attend directly to the latent unordered sequence.
A standard transformer decoder setup is used, always consisting of 6 transformer decoder units, typically with a latent space size of 64, an MLP hidden layer size of 2048, and 8 attention heads.
The unembedding is almost an exact reverse to the GMM prior embedding, acting elementwise first with a learnable MLP, then the inverse of the Cholesky-log-flatten transform described earlier for the component covariance matrix. A cross-component softmax is then applied to the unembedded weight logits to yield the GMM.
Note that for numerical stability, GMMs were always parametrized directly by the Cholesky decomposition of the covariance matrix.
For PFNs, an identical setup was used wherever possible; Observations were embedded with the same embedding architecture, and the transformer encoder used identical hyperparameters to the DT for each experiment.
Appendix C General Training Details
All experiments were trained using the Adam optimizer without weight decay or dropout (overfitting here is impossible as the model sees each sample only once). A cosine-annealing-with-warmup (5 epochs) learning rate scheduler was used. This training scheme was adopted for both DTs and PFNs.
SVI also used the Adam optimizer without weight decay or dropout, treating the parameters of the variational distribution as model parameters, and directly optimizing the ELBO via backpropagation. An exponential learning rate scheduler was used.
Experiments 4.1, 4.2.2, and 4.3 were all carried out on an NVIDIA RTX 2080 Super laptop GPU (8GB VRAM), while experiment 4.2.1 was carried out on an NVIDIA RTX 3090 GPU.
Table 4: Training hyperparameters and statistics for all experiments. An epoch always consists of 100 batches. Hyperparameters were obtained by limited manual tuning, adjusted until training was stable in all cases. For SVI, the total samples reported account for batching, so the number of samples used by each SVI problem will be a factor of the number of test problems lesser. Parameters are listed vertically, corresponding to DTs (top), PFNs (middle), and SVI (bottom). For Experiment 4.1, only DT-5 is reported but training hyperparameters are identical for DT-2. For experiment 4.3 only DT is reported. For experiment 4.2.1, two PFNs are trained (one for each dimension, as PFNs do not trivially support multivariate distributions), so two figures are quoted when necessary (PPD + hyperposterior). Model sizes are not provided for VI as this is negligible.
Experiment LR Batch Size Epochs Total Samples (M) Training Time (s) Model Size (MB)
0.005 5000 20 10 188 1.72
4.1 0.005 5000 20 10 247 1.16
0.1 10 100 10 137
0.0001 2000 200 40 4300 75.0
4.2.1 0.0001 1000 200+100 20+10 730+369 55.7
0.03 10 1000 100 131
0.001 5000 20 10 432 7.18
4.2.2 0.0001 4000 25 10 709 6.53
0.01 10 100 10 2700
4.3 0.001 5000 150 75 3720 6.86
Appendix D Details for Analytical Verification Study (Experiment 4.1)
Here, for both the inverse-gamma prior’s rate and scale parameters, we adopt an $\text{InverseGamma}(4,6)$ meta-prior and an $\text{InverseGamma}(10000,20000)$ for the wide and narrow meta-priors respectively. Both of these meta-priors have the same mean, but the narrow meta-prior is effectively singular.
An analytical expression for the ground-truth posterior distribution is obtainable via conjugacy. Specifically, given a measurement $\sigma^{2}$ and prior parameters $\alpha_{0}$ and $\beta_{0}$, the parameters of the also inverse-gamma posterior are given by:
$\alpha$ $=\alpha_{0}+\frac{1}{2},\text{ and}$
$\beta$ $=\beta_{0}+\frac{(z-\mu)^{2}}{2},$
where $z$ is the observation value and $\mu$ is the known observation mean, which we arbitrarily set to 0.
Appendix E Details for Gaussian Process Joint PPD and Hyperposterior Study (Experiment 4.2.1)
For the meta prior in this experiment, we assign a narrow, uniform prior over the PPD’s prior such that it is marginally everywhere a standard normal. We assign a uniform prior on $[1,2]$ to both parameters of the inverse-gamma prior over lengthscale.
Note that for this experiment, the PPD prior also, in conjunction with the lengthscale $l$, defines the hyperparameters of the GP’s prior mean function and RBF covariance function used here. Specifically, the GP’s mean function is set to a constant at the PPD prior’s mean, and the output scale of the RBF covariance function is set to the standard deviation of the PPD’s prior.
The sampling procedure for this experiment, using 5 observations, was as follows:
  1. 1.
    Sample $\phi$ from meta-prior
  2. 2.
    Sample $y\sim\mathcal{N}(\phi_{\mu},\phi_{\sigma^{2}})$
  3. 3.
    Sample $l\sim\text{InverseGamma}(\phi_{\alpha},\phi_{\beta})$
  4. 4.
    Sample $x\sim\text{Uniform}(0,10)$
  5. 5.
    Sample $X\sim\text{Uniform}(0,10)^{5}$
  6. 6.
    Sample $Y\sim\mathcal{N}(\phi_{\mu},k(X,X;l,\phi_{\sigma^{2}}))$
  7. 7.
    Construct $z$ as concatenation of $X$ and $Y$ in the last dimension, along with the query point $x$.
Two observation embeddings are defined, one acting on each $X$-$Y$ element pair, and one acting on the query point $x$. Each have one hidden layer of size 128.
The prior embedding consists of two MLPs, acting on the PPD prior and the lengthscale prior respectively, each with one hidden layer of size 128, with 32 outputs. These outputs are concatenated into a single vector of size 64, before being passed through another set of distinct, parallel MLPs as before to yield the required length-10 latent unordered sequence.
Appendix F Details for Quantum System Parameter Inference Study (Experiment 4.2.2)
For this experiment, we adopt a beta prior, with an $\text{InverseGamma}(4,6)$ meta-prior over each parameter.
Observations are sampled using a GPU-implemented simulation, which accepts a nominal initial state $|{\hat{\psi}_{0}}\rangle$ and nominal measurement time $\hat{t}$. We model uncertainty in the initial state preparation by, treating $|{\hat{\psi}_{0}}\rangle$ as a 2-element vector, sampling an actual initial state $|{\psi_{0}}\rangle\sim\mathcal{N}(|{\hat{\psi}_{0}}\rangle,0.01I)$, and renormalizing such that $\langle{\psi_{0}}||{\psi_{0}}\rangle=1$. We also model uncertainty in measurement time by sampling $t\sim\mathcal{N}(\hat{t},0.0025)$. We then solve Schrödinger’s equation to give $|{\psi_{t}}\rangle$, and take a measurement by sampling from the Bernoulli distribution associated with $|{\psi_{t}}\rangle$. For SVI, the likelihood of this observation is estimated stochastically with 10 samples of the initial state and measurement time.
A set of varying nominal initial states and measurement times is fixed, and with each problem sampling an observation from each is obtained. Each pair of initial state and measurement time is provided its own learnable embedding, all of standard construction. The prior embedding and posterior unembedding is also standard.
Appendix G Details for Sequential Inference Study (Experiment 4.3)
For this experiment, we use the following meta-prior over a 4-component GMM:
$w$ $\sim\text{Dirichlet}(0.25\cdot\mathbf{1_{4}})$
$x_{i}$ $\sim\mathcal{N}(\mathbf{0_{4}},4I_{4\times 4})$ $\forall i=1\ldots 4$
$\Sigma_{i}$ $\sim\text{Wishart}(5,4I_{4\times 4})$ $\forall i=1\ldots 4,$
where $\mathbf{1_{4}}$ and $\mathbf{0_{4}}$ refer to a 4-element vector consisting of only ones and zeros respectively.
We use two realistic observation models: A rangefinder subject to maximum range constraints, range-dependent Gaussian noise centred about the true range, exponentially distributed early measurements from unexpected objects, and uniformly distributed readings due to sensor failure, and bearing measurement subject to Gaussian noise and cyclic discontinuity (ie a wrapped normal distribution). Figure 7 demonstrates the complexity of the rangefinder observation model, and the parameters used in this experiment are presented in the caption.
Figure 7: Example likelihood for rangefinder observation model, conditioned on a range of 10km. We assume a maximum range of 20km, a true observation standard deviation of $0.1(\text{range}+1)$km with marginal probability 0.7, an early collision decay rate of $1\text{km}^{-1}$ with marginal probability 0.2, and a uniform sensor failure on $\text{Uniform}(0,20\text{km})$ with marginal probability 0.1. While this model is realistic, the uncertainties involved are amplified compared to those seen in reality. This was done to further illustrate the capabilities of our model.
We report iteration time per 100 series. Concretely, this refers to the prediction, update, and density fitting, of a single time step batched across 100 series.
We use a motion model representing damped velocity subject to independent, normally distributed accelerations in each direction. Formally, we adopt the following discrete-time dynamical system:
$x_{t+1}$ $=\left[\begin{matrix}1&0.63&0&0\\ 0&0.37&0&0\\ 0&0&1&0.63\\ 0&0&0&0.37\end{matrix}\right]x_{t}+\left[\begin{matrix}0.1&0.0\\ 0.2&0.0\\ 0.0&0.1\\ 0.0&0.2\end{matrix}\right]w_{t}$
$w_{t}$ $\sim\mathcal{N}(\mathbf{0_{2}},I_{2\times 2})$
$x_{0}$ $\sim\mathcal{N}(\mathbf{1_{4}},I_{4\times 4})$

### v2

@@
 We demonstrate that Distribution Transformers both maintain flexibility to vary the prior, and significantly reduces computation times—from minutes to milliseconds—while achieving log-likelihood performance on par with or superior to existing approximate inference methods across tasks such as sequential inference, quantum system parameter inference, and Gaussian Process predictive posterior inference with hyperpriors.
-Machine Learning, ICML, Approximate Inference, Transformers
 1 Introduction
-Modern machine learning has transformed many data-rich domains (LeCun et al., 2015), yet critical applications like healthcare diagnostics and autonomous vehicle safety demand more than just predictive accuracy (Hüllermeier & Waegeman, 2019).
-These high-stakes settings, often characterized by limited or noisy data, require robust uncertainty estimation and principled integration of domain knowledge (Gawlikowski et al., 2021)—capabilities naturally addressed through probabilistic reasoning.
-Bayesian inference is a cornerstone of probabilistic machine learning, encoding prior beliefs and providing the tools to incorporate new observations while providing principled uncertainty quantification; essential for critical decision-making under uncertainty.
-However, Bayesian inference is typically intractable, limiting its real-world adoption.
-To counteract this problem, the field of approximate inference has emerged, providing a range of methods: Variational inference (VI) methods (Blei et al., 2016) can approximate posterior densities but require expensive optimization procedures at inference time (as opposed to training time, relevant for later methods).
-While Markov Chain Monte Carlo (MCMC) (Christian & Casella, 2011) methods provide asymptotically-exact posterior samples, they neither directly estimate the density nor offer the rapid inference often required in practical applications.
-Recent advances in deep learning have opened new avenues for Bayesian inference.
-Prior Fitted Networks (PFNs) (Muller et al., 2021) demonstrated that transformer (Vaswani et al., 2017) architectures can perform Bayesian inference through a single forward pass, dramatically reducing inference time compared to traditional methods.
-However, PFNs are limited by their dependence on a fixed prior, requiring costly retraining for each new prior distribution encountered.
-In practice, it is often desirable, and in some applications essential, to be able to change the prior rapidly at inference time.
-While these methods have their strengths, there remains a clear need for approximate inference methods that can provide accurate posterior density estimates with minimal computational overhead while maintaining the flexibility to modify priors at inference time.
-To address this need, we introduce Distribution Transformers (DTs), a novel transformer-based architecture that combines the flexibility of VI with the efficiency of PFNs.
-Our method generates posterior density estimates through a single forward pass while handling a range of prior distributions from a given parametric family without retraining.
-Unlike traditional VI approaches that often rely on simple parametric families, DTs leverage a expressive distributional family that can universally approximate smooth densities.
-Figure 2 illustrates the architectural design of DTs.
-Fundamentally, DTs approximate the prior with a Gaussian Mixture Model (GMM) represented as an embedded sequence in the latent space of the transformer, then condition this GMM on observations using a permutation-equivariant transformer decoder, yielding a GMM approximation to the posterior.
-DTs are trained using entirely sample-based supervised learning by defining a joint distribution over the prior, parametrized by $\phi$, the latent variable $x$, and the observations $z$.
-We demonstrate that this training approach yields an approximator of exact posterior prediction.
-This combination of speed, flexibility, and expressiveness enables accurate Bayesian inference in scenarios where existing methods face fundamental limitations.
-Concretely, DTs address the key challenges discussed above, providing accurate posterior density estimates for a wide range of priors.
-Like PFNs, our approach only requires a way to sample from the chosen prior family and an arbitrary likelihood model, and is widely applicable without significant task-specific analysis.
-We make the following contributions:
-  1. 1.
-    We introduce Distribution Transformers (DTs)—a novel architecture that leverages the power of universally-approximating Gaussian Mixture Models and Transformers to express arbitrary distribution-to-distribution mappings.
-  2. 2.
-    We show that DTs can be trained to perform more accurate approximate Bayesian inference, and up to orders of magnitude faster, than existing approximate inference methods, while also allowing on-the-fly updates to the prior at inference time without retraining, a never before accomplished feat.
-  3. 3.
-    We empirically demonstrate near-exact inference capabilities on challenging problems, including joint PPD-hyperposterior approximations for Gaussian Processes (GPs), parameter inference for quantum systems and a real-world sequential inference task, while remaining fast enough for real-time applications.
+Bayesian inference provides a principled route to uncertainty quantification and the incorporation of prior knowledge.
+In practice, however, repeatedly solving inference problems (e.g., across datasets, hyperparameters, or environments) is computationally expensive.
+Amortized Bayesian inference (ABI) addresses this by learning a mapping from observations to posterior approximations, yielding fast test-time inference.
+Recent transformer-based ABI methods have shown impressive single-pass inference for small-data regimes (Muller et al., 2021; Hollmann et al., 2022, 2025).
+Yet two limitations persist.
+First, most ABI models fix the prior during training; changing the prior at test time typically requires retraining or fine-tuning.
+Second, methods that offer some prior flexibility usually do not preserve family structure between prior and posterior, hindering sequential composition (filtering/smoothing), where the posterior must become the next-step prior.
+Complementary lines in amortized SBI (Cranmer et al., 2020) and sensitivity-aware prior amortization (Elsemüller et al., 2023) underscore the need for flexible priors, but do not maintain the at-least approximate conjugacy needed for recursive updates.
+We introduce Distribution Transformers (DTs), a transformer-based architecture that (i) performs single-pass amortized inference, (ii) supports prior amortization across a family of priors at test time, and (iii) approximates conjugacy between prior and posterior, enabling clean sequential composition when needed.
+Concretely, DTs embed priors as a tokenized sequence (representing a Gaussian Mixture Model, or GMM, approximation), condition on observations with a permutation-equivariant decoder, and output a posterior in the same family.
+Training spans a distribution over priors, enabling prior-amortization.
+Concretely, we make the following contributions:
+  • A unified framework for prior-flexible, approximately conjugate amortized inference, bridging the gap between ABI and sequential inference.
+  • A practical transformer architecture that tokenizes distributions and performs posterior updates within the same parametric family.
+  • Empirical results on both static benchmarks (matching or exceeding PFN/TabPFN- and VI-style baselines) and sequential tasks (where preserving conjugacy is critical), demonstrating accuracy and speed.
 1.1 Related Work
-[…]
-Variational inference (VI) aims to approximate the posterior density through fitting a density of a particular family (typically a Gaussian) to the posterior by maximizing a variational lower bound on the evidence.
-Notable modern extensions include stochastic variational inference (SVI) (Hoffman et al., 2013), which utilizes backpropagation to optimize a stochastic approximation of the variational objective, black-box variational inference (Ranganath et al., 2014), which provides widely-applicable, model-agnostic tools for optimization of this variational objective, and amortized variational inference (AVI) (Kingma & Welling, 2014; Ganguly et al., 2023), which efficiently solves many variational inference problems in parallel through the learning of an inference network mapping from observations to approximate posteriors.
-These methods often provide acceptable estimates for the posterior density, but necessarily involve an expensive optimization routine at inference time, are not asymptotically exact, and often underestimate the posterior variance.
-Nonetheless, VI remains the most popular approximate inference technique in use today, and we will use a variant as a baseline in our empirical studies.
-Recent neural approaches to approximate inference leverage the expressivity of deep learning architectures while aiming to maintain theoretical guarantees.
-Prior Fitted Networks (PFNs) (Muller et al., 2021) learn to perform approximate inference through sample-based supervised learning without the need for posterior densities (which are instead accessed via samples from the joint distribution over $x$ and $z$), enabling inference in a single forward pass.
-They provide rapid inference for changes in dataset, provide a posterior density estimate, and allow for almost arbitrary priors, but require complete retraining upon changes to the prior distribution, and use a Riemann Distribution, which can struggle to model smooth, heavy-tailed posteriors.
-As a close competitor to our method, we will also adopt a PFN baseline for our posterior studies.
-Continuous normalizing flows (Rezende & Mohamed, 2015) learn a mapping from samples of some simple distribution to samples of the target posterior density, leveraging automatic differentiation to provide a flexible, closed-form density using a change of measure.
-However, the mapping they learn is fixed, meaning the prior is set at training time, and observations can only be incorporated via sufficient statistics of the simple distribution, limiting their suitability for our problem setting.
-Moreover, they must be constructed using only fully reversible components, which can limit expressivity.
-Classical approaches include the Laplace approximation (Tierney & Kadane, 1986) and expectation propagation (EP) (Minka, 2001).
-The Laplace approximation fits a Gaussian distribution based on the local geometry at the posterior mode, but lacks invariance to reparametrization.
-EP iteratively refines an isotropic Gaussian approximation to the posterior, but struggles with numerical stability.
-Sampling-based methods, most notably Markov chain Monte Carlo (MCMC) methods, provide samples from the posterior which can be used to compute stochastic estimates for expectations with respect to said posterior.
-These are primarily based on access to the unnormalized posterior via the numerator of Bayes’ theorem.
-When a differentiable prior and likelihood model are available, the state-of-the-art No U-Turn Sampler (NUTS) (Hoffman & Gelman, 2011) can be used.
-NUTS is based on Hamiltonian Monte Carlo (HMC) (Brooks et al., 2012), which provides more efficient generation of posterior samples without problem-specific tuning of the algorithm.
-MCMC methods are asymptotically exact and almost universally applicable, but are typically expensive due to random walk behavior.
-They also do not provide direct access to a posterior density, and empirical density estimates can be sample inefficient.
-Bayesian quadrature-based methods (Osborne et al., 2012) assign a probabilistic model, typically a GP, to the integrand of the evidence $p(z)$ (and in some cases the predictive posterior distribution PPD $p(y|x,z)=\int p(y|x)p(x\mathop{|}z)$), and use probabilistic numerical integration to estimate the evidence and PPD directly.
-These methods provide accurate uncertainty measures, and can handle noisy, black-box likelihood models, but are typically computationally expensive and scale poorly to high-dimensional problems, although recent advances in sparse-GPs (Binois & Wycoff, 2021) provide promising avenues to alleviate these scaling issues.
-1.1.1 Sequential Inference
-Sequential inference over time series requires dedicated methods, referred to as Bayesian filtering (when causality is required) or smoothing (when not).
-The extended Kalman filter (EKF) linearizes dynamics and observation models, then approximates the distribution over the latent state as a closed-form Gaussian.
-With modern automatic differentiation, this method is widely applicable and computationally efficient, but struggles in practice due to the overly restrictive assumptions induced by linearization and Gaussian distributions.
-Particle filters (PFs), or sequential Monte Carlo methods (Wills & Schön, 2023), maintain a population of particles which are distributed according to the posterior distribution each time step, but can be expensive and only provide an empirical density.
-The unscented Kalman filter (Julier & Uhlmann, 1997) maintains a minimal population of particles so is computationally efficient, but faces issues of numerical stability and also encodes a Gaussian assumption thus cannot handle multimodality.
-We compare our method against an EKF and a PF baseline.
+Amortized Bayesian Inference.
+Amortization has emerged as a key paradigm to accelerate Bayesian inference across repeated tasks by transferring expensive optimization costs to an offline training phase.
+Classical amortized variational inference (AVI) approaches (Kingma & Welling, 2013; Ganguly et al., 2023b) learn inference networks that map directly from observations to approximate posteriors, but typically rely on restrictive variational families and do not support flexible prior specification.
+Recent extensions have explored amortization in simulation-based inference (SBI), leveraging normalizing flows (Rezende & Mohamed, 2015; Papamakarios et al., 2017), neural ratio estimation (Greenberg et al., 2019; Miller et al., 2022), and hybrid MCMC-amortized approaches (Salimans et al., 2015; Gabrié et al., 2022), but these remain tied to fixed priors at training time.
+Prior-amortized methods have only recently been proposed, for example prior-amortized neural posterior estimation for reflectometry inversion (Starostin et al., 2025), and sensitivity-aware amortized inference (Elsemüller et al., 2023), highlighting both the importance and nascent state of this line of work.
+Relatedly, BayesFlow provides an amortized SBI framework based on invertible neural networks and tooling for practical workflows (Radev et al., 2020).
+Transformers for Bayesian Inference.
+Transformers have recently shown promise as amortized inference engines.
+Prior-Fitted Networks (PFNs) (Muller et al., 2021) demonstrated that transformers can approximate posterior distributions in a single forward pass, amortizing inference over datasets.
+However, PFNs assume a fixed prior, and their Riemannian output distribution struggles with smooth or heavy-tailed posteriors.
+Follow-up work extended PFNs to tabular data (TabPFN) (Hollmann et al., 2022), scaling to larger contexts and small-data regimes with improved accuracy (Hollmann et al., 2025), and to time-series forecasting (Hoo et al., 2024).
+Nonetheless, these methods remain restricted to fixed priors [^1] A rudimentary form of prior amortization can be carried out by providing prior parameters as additional observations, but this is restrictive and a side-effect not a directly-intended feature..
+In parallel, amortized in-context Bayesian inference methods (Mittal et al., 2025; Reuter et al., 2025) investigate whether transformers can learn posterior inference directly from prompts, but still lack mechanisms for prior adaptation.
+Meta-Learning and Neural Processes.
+Our work connects to neural processes (Garnelo et al., 2018a, b), which condition on context sets to predict function values, and their transformer-based extensions (Kim et al., 2019; Nguyen & Grover, 2022).
+While these frameworks amortize inference across tasks, they typically model predictive distributions over data rather than explicit posteriors over latent variables.
+ACE (Chang et al., 2024) generalized this direction to incorporate latents and flexible priors, and is thus complementary to our work.
+Other amortized meta-learning approaches (Wu et al., 2020; Iakovleva et al., 2020) have proposed shared amortized inference networks across tasks, but again without explicit prior adaptation.
+Simulation-Based Inference.
+SBI methods (Cranmer et al., 2020; Lueckmann et al., 2017; Papamakarios et al., 2019) provide amortized posterior approximations for complex simulators, with recent advances leveraging transformers and diffusion models (Gloeckler et al., 2024; Sharrock et al., 2022; Wildberger et al., 2023).
+These works are highly expressive and achieve state-of-the-art inference in scientific applications, yet typically assume a fixed prior distribution or limited parametric families.
+The lack of efficient prior amortization limits their applicability in scenarios requiring frequent prior updates, such as sensitivity analysis or robust sequential decision-making.
+Sequential Inference.
+Classical Bayesian filtering methods–such as the Kalman filter (Kalman, 1960), unscented Kalman filter (Julier & Uhlmann, 1997), and particle filters (Doucet et al., 2001; Wills & Schön, 2023)–remain the dominant approaches to sequential inference.
+While computationally efficient, they are either constrained to Gaussian assumptions or expensive particle-based representations, and provide no general amortization across tasks or priors.
+Despite the centrality of sequential inference in real-world applications, modern amortized inference frameworks (PFNs, TabPFN, ACE) have largely neglected this dimension.
+Positioning.
+In summary, prior work has established the effectiveness of amortization in approximate Bayesian inference, and the suitability of transformers for fast, in-context posterior approximation.
+However, flexible prior amortization, particularly in the context of sequential inference, remains largely unaddressed.
+Our work introduces Distribution Transformers as a principled solution, combining (i) the expressive universality of Gaussian mixtures, (ii) transformer-based amortization across priors, and (iii) applicability to sequential Bayesian filtering—providing capabilities unmatched by existing PFN, TabPFN, ACE, or simulation-based inference approaches.
 2 Preliminaries
@@
 Two theoretical properties of transformers are central to our work.
-First, transformers are universal sequence-to-sequence approximators (Yun et al., 2019), capable of representing arbitrary mappings between sequences.
-Second, in the absence of positional encodings, transformers are permutation equivariant with respect to the input sequence—a property we will exploit in Section 3.
+First, transformers are universal sequence-to-sequence approximators (Yun et al., 2019), capable of learning arbitrary mappings between sequences.
+Second, in the absence of positional encodings, transformers are permutation equivariant with respect to the input sequence—a property we exploit in Section 3.
 Our method specifically employs the transformer decoder architecture.
-This architecture extends the base transformer by incorporating global cross-attention layers, allowing each position in the input sequence to attend to a separate context sequence.
+This architecture extends the base transformer by incorporating global cross-attention layers, allowing each token in the input sequence to attend to a separate context sequence.
 This cross-attention mechanism provides a natural framework for conditioning sequence transformations on observed data.
@@
 This universality extends to distributions on compact domains under appropriate change of measure, also illustrated in Figure 1.
-While the idea of using GMMs for function approximation is not new to deep learning (for example Bishop (1994) proposes the use of a GMM to model uncertainty in the output of a neural network), the idea of operating end-to-end on a distribution represented as a GMM is novel.
+While the idea of using GMMs for function approximation is not new to deep learning (for example (Bishop, 1994) proposes the use of a GMM to model uncertainty in the output of a neural network), the idea of operating end-to-end on a distribution represented as a GMM is novel.
 Figure 1: Various log-warped GMM approximation to an inverse-gamma prior distributions.
@@
 This initial investment is offset by enabling the rapid solution of many subsequent problems, significantly reducing the overall computational burden over the model’s lifetime.
-Amortized variational inference (AVI) (Ganguly et al., 2023; Kingma & Welling, 2014) exemplifies this paradigm, where a neural network is trained to approximate posterior distributions across multiple datasets, bypassing the need for iterative optimization for each new dataset.
+Amortized variational inference (AVI) (Ganguly et al., 2023a; Kingma & Welling, 2014) exemplifies this paradigm, where a neural network is trained to approximate posterior distributions across multiple datasets, bypassing the need for iterative optimization for each new dataset.
 Similarly, PFNs (Muller et al., 2021) implicitly amortize inference through the learning of a mapping from observations to approximate posterior distributions without iterative updates.
@@
   Sample $\phi_{i}$, $x_{i}$ and $z_{i}$ from $p(\phi,x,z)$ for $i=1:b$;
-     Estimate loss ^$\hat{\ell}^{\prime}_{\theta}=-\sum_{i=1}^{b}\log q_{\theta}(f(x_{i})\mathop{|}\phi_{i})+\log q_{\theta}(f(x_{i})\mathop{|}z_{i},\phi_{i})$;
+  Estimate loss $\hat{\ell}^{\prime}_{\theta}=-\sum_{i=1}^{b}\log q_{\theta}(f(x_{i})\mathop{|}\phi_{i})+\log q_{\theta}(f(x_{i})\mathop{|}z_{i},\phi_{i})$;
   Update DT parameters with gradient descent on $\nabla_{\theta}\hat{\ell}^{\prime}_{\theta}$;
@@
 Proposition 3.1.
-The proposed loss $l_{\theta}$ equals the expected KL-Divergence $\mathbb{E}_{p(\phi,z)}\left[\text{KL}\left[p(\cdot\mathop{|}z,\phi)\mathop{||}q_{\theta}(\cdot\mathop{|}z,\phi)\right]\right]$ between $p(\cdot\mathop{|}z,\phi)$ and $q_{\theta}(\cdot\mathop{|}z,\phi)$ up to an additive constant.
+The proposed loss $l_{\theta}$ is equal to the expected KL-Divergence $\mathbb{E}_{p(\phi,z)}\left[\text{KL}\left[p\mathop{||}q_{\theta}\right]\right]$ between $p(\cdot\mathop{|}z,\phi)$ and $q_{\theta}(\cdot\mathop{|}z,\phi)$ up to an additive constant.
 Proof in Appendix A.
-The loss $l_{\theta}$ can be calculated stochastically using samples from the joint distribution $p(\phi,x,z)$, alleviating any need to directly access or sample from the posterior density.
+The loss $l_{\theta}$ can be estimated using samples from the joint distribution $p(\phi,x,z)$, alleviating any need to directly access or sample from the posterior density.
 Finally, DTs can jointly approximate priors and posteriors without additional model parameters.
 Applying the unembedding to the prior sequence yields a GMM approximation $q_{\theta}(x\mathop{|}\phi)$, extending DTs to mappings $\Phi\times\mathcal{Z}\rightarrow\Theta\times\Theta$.
-To ensure consistency, we introduce a prior loss:
+To ensure consistency and a shared latent space pre- and post-conditioning, we introduce a prior loss:
 $\ell_{\theta}^{\text{prior}}=\mathbb{E}_{p(\phi,x)}\left[-\log q_{\theta}(x\mathop{|}\phi)\right],$
-leading to the combined objective: $\ell_{\theta}^{\prime}=\ell_{\theta}^{\text{prior}}+\ell_{\theta}$
+leading to the combined objective: $\ell_{\theta}^{\prime}=\ell_{\theta}^{\text{prior}}+\ell_{\theta}$ .
 4 Empirical Studies
@@
 In the prior two settings, we benchmark our method against SVI (Hoffman et al., 2013), implemented in PyTorch (Paszke et al., 2019) with GPU parallelization, and PFNs (Muller et al., 2021), also implemented in PyTorch, fitting a Riemann distribution with the same number of model outputs as our DT.
+We share out code through an anonymised link  [^2] https://anonymous.4open.science/r/distribution-transformers-450D/.
 In the absence of a fixed prior, we train PFNs using the same sampling scheme as our method, effectively marginalizing out the meta-prior, leaving a less informative prior.
-We expect PFNs to perform well when the meta-prior is narrow and poorly when wide.
+We expect PFNs to perform well when the meta-prior is narrow and poorly when wide, as the effective prior used by the PFN is the marginalisation of the prior family with respect to the meta prior and so is close to the true prior only in the former case.
 For the latter experiment, we benchmark against the industry standard extended Kalman filter (EKF), and a particle filter (PF), again with GPU implementation.
@@
 In specific cases, where the adopted prior is of a conjugate family to the likelihood, the posterior is tractable.
-We first verify that our approach does indeed perform approximate inference, for the case of an inverse-gamma prior and normal-variance likelihood.
-We choose a meta-prior consisting of independent inverse-gamma distributions over the rate and shape parameters, and test our approach on both narrow and wide parameter settings.
+We first verify that our approach indeed performs approximate inference, for the case of an inverse-gamma prior and normal-variance likelihood.
+We choose a meta-prior consisting of independent inverse-gamma distributions over the rate and shape parameters, and test our approach on both narrow and wide settings.
 Figure 3: Ground truth, PFN, and 2 and 5 component DT posterior densities for an inverse-gamma prior with an (a) narrow and (b) wide meta-prior.
@@
 Expected KL-Divergences with the true posterior (along with 95% confidence intervals) are given for the narrow (above) and wide (below) meta-priors.
-DT-$k$ refers to a $k$-component Distribution Transformer.
+DT-$k$ refers to a $k$-component Distribution Transformer, and PFN-$n$ refers to a PFN equipped with an $n$-bucket Riemann distribution.
 Note first that both variants of our proposal achieve better posterior KL-Divergences than an equivalent PFN and SVI for both meta-prior settings, while performing inference orders of magnitude faster than SVI.
@@
 SVI
-0.0423$\pm$0.0003
-0.0558$\pm$0.0016
+[…]
+[…]
 148
-PFN
-0.814±$\pm 0.801^{*}$
-726±$\pm 1070^{*}$
-0.0144
+PFN-15
+[…]
+[…]
+0.003
+PFN-5000
+[…]
+[…]
+0.003
+TabPFNv2
+[…]
+[…]
+1.52
+ACE-5
+[…]
+[…]
+0.037
 DT-2
-0.0045$\pm$0.0001
-0.0058$\pm$0.0002
-0.0150
+$4±\pm0.00$01
+[…]
+0.014
 DT-5
-0.0004$\pm$0.0000
-0.0003$\pm$0.0000
-0.0156
+[…]
+[…]
+0.016
 Figure 3 demonstrates that for both wide and narrow meta-priors, the Distribution Transformer does indeed learn to perform Bayesian inference and provides excellent approximations to the posterior, even under change of prior.
 Figure 1 shows the GMM approximations provided by the DT for this study, demonstrating that even with no additional model parameters a high-quality mapping to a GMM approximation of the prior is achieved.
-It is clear that while PFNs perform as well as the Riemann distribution allows for the fixed-prior case, as expected in the case of variable priors they fail to fit the posterior at all.
+It is clear that while PFNs perform as well as the Riemann distribution allows for the narrow meta-prior case, as expected in the wide case they fail to fit the posterior at all.
 This is further demonstrated in Table 1, which shows that not only do we achieve speeds close to that of PFNs, which are slightly faster due to the lighter-weight architecture, but even significantly outperform the much slower SVI in terms of posterior KL-Divergence.
 These performance gains can be attributed to the expressive GMM adopted by our method, and the efficient transformer-based architecture.
+TabPFNv2, while being much bigger of a model than any other baseline, is also unable to improve much on top of standard PFNs.
+ACE achieves second best performance, losing only to DTs.
+It thus appears that all methods utilising GMMs surpass those based on Riemannian histogram.
+However, DTs can provide greater improvement, presumably because self-attention is applied to mixture components (as opposed to data in ACE) allowing the model to spend more computations on shaping the correct distribution shape.
 An interesting observation, is the extremely high uncertainty in the estimate for the PFN posterior KL-Divergence, marked *.
@@
 We now demonstrate that our method can quickly perform approximate inference jointly over both the predictive posterior and hyperposterior.
-In Table 2 we show that we outperform existing methods in terms of NLL for both PPD and hyperposterior, as expected, while also being the fastest method in terms of runtime.
-In Figure 4, we show an example PPD for for one set of datapoints and are clearly closer to the oracle PPD that has the knowledge of the true lengthscale.
-Figure 4: Example plot for the GP predictive experiment with hyperpriors, using 10 GMM components.
+In Table 2 we show that we outperform existing methods on a more challenging 5-dimensional input problem in terms of NLL for both PPD and hyperposterior, as expected, while also being the fastest method in terms of runtime.
+In Figure 4, we show an example PPD for for one set of datapoints and are clearly closer to the oracle PPD knowing the true lengthscale.
+Figure 4: Example plot for the 1-dimensional input GP predictive experiment with hyperpriors, using 10 GMM components.
 Here we put an InverseGamma(1,2) prior on the lengthscale.
@@
 In orange, we show the oracle that is the exact GP fit with the true lengthscale value (which is unobserved for the other methods).
-We see that PFNs overestimate the confidence intervals due to the fact that they do not take the prior, particularly that of the lengthscale, into account.The Riemann distribution of the PFN uses 30 buckets.
+We see that PFNs overestimate the confidence intervals due to the fact that they do not take the prior, particularly that of the lengthscale, into account.The Riemann distribution of the PFN uses 30 buckets, matching the number of outputs of the DT.
 Table 2: Expected NLL for the marginal PPD and the marginal hyperposterior (denoted Hyper), and inference time per 1000 problems (denoted Runtime), for Experiment 4.2.1.
@@
 As expected, we outperform in both categories.
-Note that PFNs suffer the same issue with the hyperposterior NLL confidence here as the one in Experiment 4.1.
-Method Expected NLL
-PPD Hyper Runtime (s)
-VI 1.97 $\pm$ 0.53 131
-PFN 0.00 $\pm$ 0.10 3.21 $\pm$ 4.40 1.03
-DT -0.32 $\pm$ 0.12 0.81 $\pm$ 0.08 0.96
+Note that PFNs suffer the same issue with the hyperposterior NLL confidence here as in Experiment 4.1.
+The usually-accurate MCMC also underperforms here, as certain meta-prior settings give rise to instability and non-convergence, and thus inaccuracies.
+Method Expected NLL Inference
+PPD Hyper Time (s)
+VI 0.39 $\pm$ 0.05 123
+MCMC 2.27 $\pm$ 0.04 27
+PFN 0.90 $\pm$ 0.02 1.53 $\pm$ 0.01 9.2
+tabPFNv2 1.08 $\pm$ 0.02 0.37 $\pm$ 0.02 28.5
+ACE 1.08 $\pm$ 0.02 0.35 $\pm$ 0.02 11.7
+DT 0.81 $\pm$ 0.02 0.31 $\pm$ 0.02 9.5
 4.2.2 Quantum System Parameter Inference
@@
 We model observations as 10 independent experiment runs, subject to uncertainty in initial state preparation and measurement times, and generated via a GPU-implemented simulation.
-Figure 5: Expected negative log-likelihood against batch inference time for SVI, a PFN and a 5-component DT.
+Figure 5: Expected negative log-likelihood against batch inference time.
 Note that even given orders of magnitude more computation time, SVI cannot match the performance of our method, again demonstrating the power of our GMM approximation.
 Note that this problem is particularly challenging for VI, as the likelihood must be marginalized with respect to the uncertainty in the initial state and measurement time, which is not tractable and must be estimated stochastically, increasing the time per iteration.
-Figure 5 illustrates the power of our approach, achieving better log-likelihood performance (and therefore a smaller KL-Divergence with the true posterior, indicating a better fit), than both competitor methods.
-This is a problem setting where variable priors are important, as they can be subjective in this setting, and prior sensitivity analysis can be important.
-To enable such an analysis, fast inference is also important.
-This illustrates the practical advantage of our approach over PFNs and VI well, as we explicitly target these problems.
+Figure 5 illustrates the power of our approach, achieving better log-likelihood performance (and therefore a smaller KL-Divergence with the true posterior, indicating a better fit), than PFNs and SVI and matching performance of TabPFNv2 and ACE, while being faster.
+This is a problem setting where variable priors are important, as the prior sensitivity analysis is typically necessary.
+To enable such an analysis, fast inference is crucial.
+This further illustrates the practical advantage of our approach, as we explicitly target these problems.
 4.3 Sequential Inference Study
@@
 However, these assumptions are rarely reflected in reality.
-We tackle a sensor fusion problem consisting of 2-dimensional linear dynamics, modelled as a 4-dimensional state space (displacement and velocity), with indirect measurements of displacement provided by two independent, non-linear, non-Gaussian, sensors.
-Our approach is likely to be successful here, as prior time spent training is almost irrelevant, and fast inference with variable priors is the priority.
+We tackle a sensor fusion problem consisting of 2-dimensional linear dynamics, modelled as a 4-dimensional state space, with indirect measurements of displacement provided by two independent, non-linear, non-Gaussian, sensors.
+Our approach is well-suited to this problem setting, as time spent training is almost irrelevant, and fast, conjugate inference with variable priors is the priority.
 We also validate against a particle filter, which handles these non-ideal conditions well at the expense of computation time.
@@
 Note that the DT generally has a higher uncertainty than the EKF, demonstrating proper handling of the complex uncertainty structure of the observations.
-Table 3: Expected NLL and iteration time (prediction, update and distribution generation) for EKF, PF, and DT.
-Note that our method vastly outperforms the EKF in terms of NLL with a small cost in iteration time, while almost matches the close to ground-truth PF’s NLL while achieving close to 50$\times$ speedup.
+[…]
 Method Expected NLL
@@
 100 Series Batch (s)
-EKF 95.9$\pm$4.40 0.010
-PF -0.244$\pm$0.047 0.818
-DT -0.197$\pm$0.040 0.017
+[…]
+[…]
+[…]
 Figure 6 clearly demonstrates that our approach tracks the true state well, while the EKF fails to track the true state at all.
@@
 5 Conclusion
-In summary, Distribution Transformers (DTs) introduce a powerful framework for approximate Bayesian inference by combining universally-approximating Gaussian Mixture Models with Transformer architectures.
-Our empirical results demonstrate that DTs not only achieve superior inference accuracy and speed compared to existing methods, but also enable dynamic prior updates without retraining—a unique capability in the field.
-Furthermore, DTs show near-exact inference performance on challenging tasks including GP hyperposterior estimation, quantum parameter inference, and real-world sequential inference, while maintaining the computational efficiency needed for practical applications.
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+In summary, Distribution Transformers (DTs) introduce a powerful framework for approximate Bayesian inference by combining universally-approximating GMMs with Transformer architectures.
+Our empirical results demonstrate that DTs not only achieve superior inference accuracy and speed compared to existing methods, but also enable dynamic prior updates without retraining while maintaining conjugacy—a unique capability in the field.
+Furthermore, DTs show competitive inference performance in terms of expected posterior NLL on challenging tasks, while maintaining the computational efficiency needed for practical applications.
+Checklist
+  1. 1.
+    For all models and algorithms presented, check if you include:
+    1. (a)
+      A clear description of the mathematical setting, assumptions, algorithm, and/or model. Yes
+    2. (b)
+      An analysis of the properties and complexity (time, space, sample size) of any algorithm. Not Applicable
+    3. (c)
+      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. Yes
+  2. 2.
+    For any theoretical claim, check if you include:
+    1. (a)
+      Statements of the full set of assumptions of all theoretical results. Yes
+    2. (b)
+      Complete proofs of all theoretical results. Yes
+    3. (c)
+      Clear explanations of any assumptions. Yes
+  3. 3.
+    For all figures and tables that present empirical results, check if you include:
+    1. (a)
+      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). Yes
+    2. (b)
+      All the training details (e.g., data splits, hyperparameters, how they were chosen). Yes
+    3. (c)
+      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). Yes
+    4. (d)
+      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). Yes
+  4. 4.
+    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
+    1. (a)
+      Citations of the creator If your work uses existing assets. Not Applicable
+    2. (b)
+      The license information of the assets, if applicable. Not Applicable
+    3. (c)
+      New assets either in the supplemental material or as a URL, if applicable. /Not Applicable
+    4. (d)
+      Information about consent from data providers/curators. Not Applicable
+    5. (e)
+      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. Not Applicable
+  5. 5.
+    If you used crowdsourcing or conducted research with human subjects, check if you include:
+    1. (a)
+      The full text of instructions given to participants and screenshots. Not Applicable
+    2. (b)
+      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. Not Applicable
+    3. (c)
+      The estimated hourly wage paid to participants and the total amount spent on participant compensation. Not Applicable
+Supplementary Materials
 Appendix A Proof of Proposition 3.1
@@
 A standard transformer decoder setup is used, always consisting of 6 transformer decoder units, typically with a latent space size of 64, an MLP hidden layer size of 2048, and 8 attention heads.
+Layer norm was also chosen for its numerous benefits (Ziomek et al., 2025).
 The unembedding is almost an exact reverse to the GMM prior embedding, acting elementwise first with a learnable MLP, then the inverse of the Cholesky-log-flatten transform described earlier for the component covariance matrix.
@@
 For PFNs, an identical setup was used wherever possible; Observations were embedded with the same embedding architecture, and the transformer encoder used identical hyperparameters to the DT for each experiment.
+For ACE, we also tried to use an similar setup as possible; for this region we used the GMM heads and used the same number of components and sample space transformations as in DTs.
+Additionally, since we are only interested in inference of certain variables in the problem, we did not randomise the variable with respect to which NLL loss is computed and instead kept it fixed to the variable of interest.
+We provide prior hyperparameters as additional latent variables on each problem.
+For TabPFNv2, we simply used the pretrained foundational model provided by the authors and conditioned it on 10,000 samples (which is the maximum recommended by the authors).
+We provided prior hyperparameters as additional observations the model was conditioned on.
 Appendix C General Training Details
@@
 For SVI, the total samples reported account for batching, so the number of samples used by each SVI problem will be a factor of the number of test problems lesser.
-Parameters are listed vertically, corresponding to DTs (top), PFNs (middle), and SVI (bottom).
+Parameters are listed vertically, corresponding to DTs (top), PFNs (second from top), SVI (third from top) and ACE (bottom).
 For Experiment 4.1, only DT-5 is reported but training hyperparameters are identical for DT-2.
@@
 0.1 10 100 10 137
+0.001 5000 20 10 171 3.93
 0.0001 2000 200 40 4300 75.0
@@
 0.03 10 1000 100 131
-0.001 5000 20 10 432 7.18
-4.2.2 0.0001 4000 25 10 709 6.53
-0.01 10 100 10 2700
+0.0001 5000 40 80 4160 95.8
+0.001 5000 20 10 222 7.18
+4.2.2 0.0001 4000 25 10 290 6.53
+0.01 10 100 10 1036
+0.0001 5000 20 10 356 9.22
 4.3 0.001 5000 150 75 3720 6.86
@@
 For the meta prior in this experiment, we assign a narrow, uniform prior over the PPD’s prior such that it is marginally everywhere a standard normal.
-We assign a uniform prior on $[1,2]$ to both parameters of the inverse-gamma prior over lengthscale.
+We assign an inverse-gamma prior over lengthscale, whose hyperparameters are sampled from a uniform meta-prior, which is over $[100,105]$ for concentration and $[100,700]$ for rate.
 Note that for this experiment, the PPD prior also, in conjunction with the lengthscale $l$, defines the hyperparameters of the GP’s prior mean function and RBF covariance function used here.
 Specifically, the GP’s mean function is set to a constant at the PPD prior’s mean, and the output scale of the RBF covariance function is set to the standard deviation of the PPD’s prior.
-The sampling procedure for this experiment, using 5 observations, was as follows:
+The sampling procedure for this experiment, using 5 observations (each observation being 5-dim input and 1-dim function value), was as follows:
   1. 1.
@@
   4. 4.
-    Sample $x\sim\text{Uniform}(0,10)$
+    Sample $x\sim\text{Uniform}(0,5)^{5}$
   5. 5.
-    Sample $X\sim\text{Uniform}(0,10)^{5}$
+    Sample $X\sim\text{Uniform}(0,5)^{5\times 5}$
   6. 6.
@@
 Appendix F Details for Quantum System Parameter Inference Study (Experiment 4.2.2)
-For this experiment, we adopt a beta prior, with an $\text{InverseGamma}(4,6)$ meta-prior over each parameter.
+This experiment is a probabilistic adaptation of a well-known two-level quantum system problem (Nielsen & Chuang, 2010; Schorling et al., 2025).
+We adopt a beta prior, with an $\text{InverseGamma}(4,6)$ meta-prior over each parameter.
 Observations are sampled using a GPU-implemented simulation, which accepts a nominal initial state $|{\hat{\psi}_{0}}\rangle$ and nominal measurement time $\hat{t}$.